OpenStudy (anonymous):

£400 is divided between Ann, Brett and Carol, so that Ann has twice as much as Brett, who has three times has much as Carol, How much does Brett receive???

OpenStudy (anonymous):

(2*3)A + 3B + C = 400

OpenStudy (anonymous):

i do not get that.

OpenStudy (anonymous):

What do you get? Think of it. If Brett has 3 times as much as Carol that means C*3=B, and Anna has twice as much as Brett, then A=2B, so you have 6A + 3B + A = 400

OpenStudy (anonymous):

I mean 6A + 3B + C = 400, then go to total is 10, so 400/10 = 40. So what is 6*40 + 3*40 + 40?

OpenStudy (anonymous):

actually I am confused

OpenStudy (anonymous):

where did you get the 6a+3b+c

OpenStudy (anonymous):

Ok you here?

OpenStudy (anonymous):

yessir

OpenStudy (anonymous):

Sounds good. I got 3b because brett has 3 times as much as carol. think of it as two people at first. then i got 6a cause anna has 2 times as much as brett. so you would take 2*3*a=6a which is what anna has and 3*b=3b and then carol is just left with her even amount. So you get 6a+3b+c=400..notice how if you add 6+3+1 you get 10...then you can take 400/10 to get 40. and then multiply 40 back into 6a and 3b and c to get 240, 120, and 40. A check is to put the answer back in. when you do you get 400. You can also think of it as 10 different people getting 400. So each person get 40. The hardest part here is noticing the 10. You just have to think that 3b is how much c has and 2*3 is how much a has and c is left with just itself. This is really long sorry if it's hard to understand. Let me know if I can make it any more clearer. Good luck understanding it!

OpenStudy (anonymous):

FANTASTIC